# gamecenter,boxed

### v1

@@ 3. Problem Statement
$\leq$ $\exp(\epsilon)\cdot\Pr[\mathcal{A}(\textsc{view}_{A}^{\Pi}(D_{A},D^{\prime}))=1]+negl(\lambda).$
@@
Intuitively, after observing the output and intermediate computations, the amount of information that any party can learn about any other party’s local dataset should be bounded by differential privacy.
@@
  • The existence of an honest-but-curious service provider S to assist with multiparty computations. This party should be oblivious to all inputs and results, i.e., the service provider learns nothing about the client’s input or output during the execution of the protocol. We discuss alternatives to this assumption in Section 7.
  • Clients share a common secret before the start of the protocol; this can be used as a seed to a common Pseudo-Random Number Generator (PRNG) and is generally possible through public-key schemes or communication over a channel secured by TLS.
4. Related Work
4.1. Secure Exact $k$-means
We first discuss related work using secure computation to compute $k$-means that does not preserve the privacy of the output. This literature shows a trade-off between privacy and efficiency. Approaches that prioritize efficiency often compromise privacy by allowing the leakage of intermediate computations such as sums, counts, or centroids, which can be exploited to infer sensitive data (Gheid and Challal, 2016; Jiang et al., 2020). For instance, Gheid et al. (Gheid and Challal, 2016) rely on computing secure sums for aggregating sums and counts across clients. While this is very efficient in practice, it reveals aggregate intermediate values to all clients, which is not considered secure (Hegde et al., [n. d.]). Another approach by Jiang et al. (Jiang et al., 2020) employs homomorphic encryption and garbled circuits to conceal sums and counts. However, it still reveals the intermediate centroids to all clients in the clear. Many other works in this category are explained in the SoK of Hedge et al. (Hegde et al., [n. d.]).
Conversely, secure approaches that do not leak intermediate computations employ heavy cryptographic techniques, like fully homomorphic encryption, leading to significant performance degradation (Bunn and Ostrovsky, 2007; Rao et al., 2015; Jäschke and Armknecht, 2018; Kim and Chang, 2018). Bunn and Ostrovsky (Bunn and Ostrovsky, 2007) utilize arithmetic secret sharing alongside homomorphic encryption, with high-performance overheads. Homomorphic encryption has also been used by Rao et al. (Rao et al., 2015), Jäschke et al. (Jäschke and Armknecht, 2018), and Kim et al. (Kim and Chang, 2018), and consistently results in prohibitively high computation time due to the computation demands of the encryption. The work of Mohassel et al. (Mohassel et al., 2020) is a notable exception, offering a scheme that is significantly faster than previous state-of-the-art methods. By using optimized $2$-party computation primitives, batched oblivious transfers, and garbled circuits, they achieve a speedup of five orders of magnitude over Jäschke et al. (Jäschke and Armknecht, 2018). However, the runtime is still on the order of minutes; and the communication size is on the order of gigabytes. In contrast, our approach is as efficient as the protocols that leak intermediate computations, while protecting the privacy of the output and intermediate computations with DP.
4.2. Differentially Private $k$-means
@@
The central model operates with a trusted curator who collects data for analysis, hiding individuals’ information in the output, but not from the curator itself. Among the central approaches, DP-Lloyd introduces noise during each centroid update, by adding Laplace noise to both the sum (numerator) and the counts (denominator) when computing the arithmetic mean of points within a cluster (Blum et al., 2005). DP-Lloyd was later implemented in the PinQ framework (McSherry, 2009) with a fixed number of iterations. Dwork (Dwork, 2011) extended this framework to allow for an arbitrary number of iterations, allocating an exponentially decreasing privacy budget to each iteration, although it was noted that utility degrades beyond a certain number of iterations due to increasing noise levels. The state-of-the-art in DP-Lloyd was achieved by Su et al. (Su et al., 2016), who included an error analysis to determine the optimal number of iterations and introduced a “sphere-packing” centroid initialization method that significantly improves clustering quality.
@@
The Sample and Aggregate Framework (SaF) partitions the dataset into multiple subsets, upon which the non-private Lloyd’s algorithm is executed independently (Nissim et al., 2007). The resulting centroids from each subset are then aggregated using a standard DP mechanism. SaF was later implemented within the GUPT system (Mohan et al., 2012). However, experiments conducted by Su et al. (Su et al., 2016) suggest that DP-Lloyd consistently outperforms SaF across a variety of synthetic and real datasets. Synopsis-based methods take a different approach by first generating a synopsis of the dataset using a differentially private algorithm, then applying the $k$-means clustering algorithm on this synopsis. Qardaji et al. (Qardaji et al., 2012) proposed a synopsis technique for 2D datasets, which was extended and optimized by Su et al. (Su et al., 2016) to higher dimensions. However, this approach only outperforms DP-Lloyd in datasets with less than three dimensions (Su et al., 2016). Another, more theoretical, line of work focuses on minimizing approximation error bounds, without implementation or empirical evaluation (Feldman et al., 2009, 2017; Balcan et al., 2017; Nissim and Stemmer, 2018; Stemmer and Kaplan, 2018; Ghazi et al., 2020). The work of Ghazi et al. (Ghazi et al., 2020) concludes this long line of work by achieving the same approximation factor as the best non-private algorithms. However, these methods suffer from superlinear running times, making them impractical for large datasets, where privacy is most important.
Local DP
@@
In contrast to central DP, the local model operates under the premise that no trusted curator is available. This necessitates that individuals randomize their data locally before it is aggregated, which reduces utility by introducing a significantly higher level of noise (on the order of the number of clients). Most research in this area, including works by Nissim and Stemmer (Nissim and Stemmer, 2018) and Stemmer (Stemmer, 2021), focus on theoretical error bounds rather than practical applications. Finally, Li et al. (Li et al., 2023) attempt to bridge the gap between these central and local models in the vertically federated setting. Their approach relies on an untrusted curator aggregating noisy local centers and membership encodings from the clients to generate a private synopsis. However, this approach is specifically tailored for the vertical setting and cannot be applied to our configuration. Our work yields utility that is even better than state-of-the-art work in the central model (Su et al., 2016) while obtaining the trust assumptions of the local model and operating in the federated setting.
5. Distributed DP $k$-means
@@
First, we overview our modified DP algorithm, using constrained $k$-means. Then, we describe FastLloyd, including the initialization, assignment, and update steps. Finally, we provide an error analysis and proof of privacy.
5.1. Constrained DP $k$-means
We observe that neither federated exact $k$-means nor DP $k$-means algorithms in the literature take advantage of constrained clustering approaches from the non-private literature (Bennett et al., 2000; Levy-Kramer, 2018). In this section, we describe how incorporating constrained clustering into the assignment step significantly improves utility; a contribution that is independently applicable to the central model of DP.
In the DP $k$-means literature (Blum et al., 2005; McSherry, 2009; Dwork, 2011; Su et al., 2016), the typical way to compute a differentially-private mean is to perform two differentially-private queries: one for the sum and one for the count. The sum is then divided by the count as a postprocessing step. We call this the sum-count method. If the datapoints $x_{l}$ are contained in $[-B,B]$, then the sensitivity of the sum is $B$, and the sensitivity of the count is $1$. Adding noise at these sensitivities yields reasonable utility, since the sum has an expected magnitude much larger than $B$, so the noise does not dominate the value, and upon division, the noise is further reduced.
The sum-count method is typically more privacy-efficient than releasing the mean directly. If we release the mean directly, the sensitivity needs to account for the existence of clusters with a single data point, leading to a sensitivity of $B$ for the mean. Since the magnitude of the mean itself is at most $B$, a sensitivity of $B$ produces a relatively large amount of noise, leading to unusable utility.
We show that by using constrained clustering, we can tighten the sensitivity bound and release the mean directly. This avoids dividing the privacy budget between the sum and count queries and gives a much higher utility than the sum-count method as we show in Section 6.2. By restricting the minimum cluster size to $\eta_{min}$, we provide a tunable bound on the sensitivity of the mean. Since all clusters have at least $\eta_{min}$ data points, the denominator of the centroid is now bounded below by $\eta_{min}$. Therefore, a single data point cannot affect any centroid by more than $\frac{B}{\eta_{min}}$.
(6) $\Delta^{\mu}=\frac{B}{\eta_{min}}$
The centroid-method for computing the DP mean is:
(7) $\bar{\mu_{j}}=\frac{\sum_{x_{l}\in O_{j}}x_{l}}{|O_{j}|}+\gamma^{\mu}\quad\text{where}\quad\gamma^{\mu}\sim Lap\left(\frac{\Delta^{\mu}}{\epsilon^{\mu}}\right)$
We further claim that, by employing constrained clustering, the sum-count method can also be improved. Intuitively, having a minimum cluster size prevents the formation of very small clusters, where the differentially-private noise would dominate the underlying sum and count. In Section 6.2, we show that the sum-count method is improved by using constrained clustering, but not as much as the centroid method.
5.2. Overview of FastLloyd
In this section, we describe our protocol for federated DP $k$-means, FastLloyd. Algorithm 1 overviews the protocol from the perspective of a single client. Following the structure of Lloyd’s algorithm, the following sections describe how we design each of its main steps: initialization, assignment, and update.
5.2.1. Initialization: Sphere Packing
@@
We modify the initialization so that it can be carried out in a federated manner. We employ the sphere packing initialization approach of Su et al. (Su et al., 2016) as it was shown to outperform random initialization. The sphere packing approach is data-independent and thus does not use any privacy budget. The process can be outlined as follows:
  1. (1)
@@
  3. (3)
    If, after repeated attempts, it is not possible to find such a point, decrease the radius $a$ and repeat the process.
@@
5.2.2. Assignment: Constrained $k$-means
To set the minimum cluster ($\eta_{min}$) and maximum cluster ($\eta_{max}$) sizes, we apply a heuristic of setting them as a ratio of the average cluster as follows:
(8) $\eta_{min}=\frac{N}{\alpha_{min}k}\quad\text{and}\quad\eta_{max}=\frac{\alpha_{max}N}{k}$
where $\alpha_{min}$ and $\alpha_{max}$ are hyperparameters that control the minimum and maximum cluster sizes. In Appendix B, we show that $\alpha_{min}=1.25$ and $\alpha_{max}=1.25$ provide a good balance between utility and feasibility. We use the open-source implementation from Levy-Kramer (Levy-Kramer, 2018) to enforce the constraints in the assignment step. In Line 6 of Algorithm 1, the client computes their assignments locally using the min-cost flow solver.
@@ 5.2.2. Assignment: Constrained $k$-means
2:Output: Cluster Centres $\boldsymbol{\mu}^{(T)}$
3:$\boldsymbol{\mu}^{(0)}=$ Initialization($seed$)
@@
6:     Compute $\mathbf{O}_{i}^{(t)}=$ MinCostFlowSolve($D_{i},\boldsymbol{\mu}^{(t-1)}$)
7:     // Local Update
8:     for $j\in\{1,\dots,k\}$ do
9:         Compute $\bar{\mu}_{ij}^{(t)}=\frac{\sum_{x_{l}\in O_{ij}^{(t)}}x_{l}}{|O_{ij}^{(t)}|}$
10:     // Global Update
11:     $\widehat{\boldsymbol{\mu}}=$ GlobalMSA($\mathbf{\bar{\mu}}_{i}^{(t)}/M$, $\Delta$, $\epsilon^{\mu},seed,t$)
12:     // Post Process Result
13:     $\boldsymbol{\mu}^{(t)}=$ Fold($\widehat{\boldsymbol{\mu}}$)
5.2.3. Local Update
The client has already locally assigned each of their data points to a cluster, following the constraints, in the previous step (Line 6). Next, they compute the sum divided by the count of each cluster, using their local dataset. We call this the local update, and it follows the classic Lloyd’s algorithm exactly. The output of this step is a matrix (of dimensions $(k,d)$) containing the average of each dimension for each cluster on client $p_{i}$’s data.
5.2.4. Global Update: Masked Secure Aggregation
We modify the update step to release a perturbed version of the cluster centres to each client under the CDP security model. By publishing the previous noisy centroids, the assignment and local centroid computation can be performed locally by each client instead of using a secure computation protocol. In the global update step (Line 11), we privately aggregate (average) the local centroids of all clients, revealing the noisy global centroids for the next round. The clients scale their local centroids down by $M$, before calling the global update so that the sum of the local updates results in the average of the local centroids.
We describe our aggregation protocol in Figure 2. We call this protocol Masked Secure Aggregation (MSA). In MSA, similar to regular secure aggregation protocols (Bonawitz et al., 2017), a service provider S securely aggregates values from $M$ clients, all while oblivious to every client’s contributions. However, in MSA, the server is also oblivious to the result of the aggregation operation; the server only acts as an aggregator (who also adds DP noise) and is oblivious to both the input and output. Each client $p_{i}$ possesses a private value (a matrix $v_{i}\in\mathbf{R}^{k\times d}$, where $\mathbf{R}$ could be $\mathbb{Z}_{2^{64}}$). These clients aim to collectively compute the element-wise sum of their private matrices: $\boldsymbol{v}=\sum_{i=1}^{M}{\frac{v_{i}}{M}}$. In addition to the client’s values, we assume the privacy parameters ($\epsilon,\Delta$) and a shared random $seed$ are also provided.
Clients Send Data
The first step in the protocol is for each of the $M$ clients to generate a random mask set. The random mask set is composed of $M$ random matrices of the same dimensions as the client items: $\{r_{1},r_{2},\dots r_{M}\}$ where each value is sampled uniformly from $\mathbf{R}$. Because all the clients have access to a shared seed and $PRNG$ they can each compute the entire set locally. In Line 2, each client samples the entire set and sums it (in Line 2) to get the global mask matrix $\boldsymbol{r}$ (which will be used to decrypt later).
In Line 2, each private value is converted to a fixed-point format, $\tilde{v_{i}}$, by scaling it up with a power-of-2 scale factor, $\mathbf{2^{\delta}}$. Formally, $\tilde{v_{i}}=\lceil{v_{i}\times\mathbf{2^{\delta}}}\rfloor$. All fractional values are rounded to the nearest representable value in this fixed-point representation. The choice of the scale factor, $\mathbf{2^{\delta}}$, determines the precision of the representation [^1] A larger $\mathbf{2^{\delta}}$ allows for greater precision but also reduces the number of bits available for the integer part of the number, which might cause overflows. A workaround is to increase the bit-width of the operations, which increases the computational load and the communication cost.. In practice, we empirically choose $\delta=16$.
Then, in Line 2, each client encrypts (masks) their input $\tilde{v_{i}}$ by adding their share of the mask $r_{i}$. Finally, each client sends their masked values to the server. This step masks the actual value $v_{i}$ from the server because $r_{i}$ acts as a one-time pad. Note that this masked value is never sent to other clients, as they would be able to unmask it easily.

### v2

@@ 3 Problem Statement
$\leq$ $\exp(\epsilon)\cdot\Pr[\mathcal{A}(\textsc{view}_{A}^{\Pi}(D_{A},D^{\prime}))=1]+negl(\lambda)+\delta.$
@@
Intuitively, after observing the output and intermediate computations, the amount of information that any party can learn about any other party’s local dataset should be bounded by differential privacy. Compared to the local and shuffle models of DP, the IND-CDP-MPC model provides a significantly better privacy vs. utility trade-off at the level of the central model of DP. We show that removing the need for a trusted aggregator (needed in the central DP model) does not require the computational resources of a large-scale secure computation protocol. We achieve this by processing data in encrypted form (secure computation), accelerated by revealing differentially private intermediate values.
@@
  • The existence of an honest-but-curious service provider S to assist with multiparty computations. This party should be oblivious to all inputs and results, i.e., the service provider learns nothing about the client’s input or output during the execution of the protocol (a weaker assumption than the local and shuffle models). We also discuss alternatives to this assumption in Appendix C.
  • Clients share a common secret used as a PRNG seed and do not collude with the server. This secret can be established through standard public-key schemes (e.g., Bonawitz et al. [10] uses Diffie-Hellman key exchange [14]) or TLS-secured channels.
4 Related Work
4.1 Secure Exact $k$-means
We first discuss secure computation based approaches to $k$-means that do not preserve the privacy of the output. This literature shows a trade-off between privacy and efficiency. Approaches that prioritize efficiency often compromise privacy by allowing the leakage of intermediate computations such as sums, counts, or centroids, which can be exploited to infer sensitive data [29, 38]. For instance, Gheid et al. [29] rely on computing secure sums for aggregating sums and counts across clients. While this is very efficient in practice, it reveals aggregate intermediate values to all clients, which is not considered secure [31]. As we discuss in Appendix C, naively fixing this leakage would require a significant degradation in utility or performance. Another approach by Jiang et al. [38] employs homomorphic encryption and garbled circuits to conceal sums and counts. However, it still reveals the intermediate centroids to all clients in the clear. Many other works in this category are explained in the SoK of Hedge et al. [31].
Conversely, secure approaches that do not leak intermediate computations employ heavy cryptographic techniques, like fully homomorphic encryption, leading to significant performance degradation [11, 59, 36, 40]. Bunn and Ostrovsky [11] utilize arithmetic secret sharing alongside homomorphic encryption, with high-performance overheads. Homomorphic encryption has also been used by Rao et al. [59], Jäschke et al. [36], and Kim et al. [40], and consistently results in prohibitively high computation time due to the computation demands of the encryption. The work of Mohassel et al. [50] is a notable exception, offering a scheme that is significantly faster than previous state-of-the-art methods. By using optimized $2$-party computation primitives, batched oblivious transfers, and garbled circuits, they achieve a speedup of five orders of magnitude over Jäschke et al. [36]. However, the runtime is still on the order of minutes; and the communication size is on the order of gigabytes. In contrast, our approach is as efficient as the protocols that leak intermediate computations, while protecting the privacy of the output and intermediate computations with DP.
4.2 Differentially Private $k$-means
@@ 4.2.1 Central DP
The central model operates with a trusted curator who collects data for analysis, hiding individuals’ information in the output, but not from the curator itself. Among the central approaches, DP-Lloyd introduces noise during each centroid update, by adding Laplace noise to both the sum (numerator) and the counts (denominator) when computing the arithmetic mean of points within a cluster [9]. DP-Lloyd was later implemented in the PinQ framework [47] with a fixed number of iterations. Dwork [16] extended this framework to allow for an arbitrary number of iterations, allocating an exponentially decreasing privacy budget to each iteration, although it was noted that utility degrades beyond a certain number of iterations due to increasing noise levels. The state-of-the-art in DP-Lloyd was achieved by Su et al. [64], who included an error analysis to determine the optimal number of iterations, optimized the splitting of privacy budget between sum and count queries, and introduced a “sphere-packing" centroid initialization method that significantly improves clustering quality.
@@
The Sample and Aggregate Framework (SaF) partitions the dataset into multiple subsets, upon which the non-private Lloyd’s algorithm is executed independently [53]. The resulting centroids from each subset are then aggregated using a standard DP mechanism. SaF was later implemented within the GUPT system [49]. However, experiments conducted by Su et al. [64] suggest that DP-Lloyd consistently outperforms SaF across various synthetic and real datasets. Synopsis-based methods take a different approach by first generating a synopsis of the dataset using a differentially private algorithm, then applying the $k$-means clustering algorithm on this synopsis. Qardaji et al. [57] proposed a synopsis technique for 2D datasets, which was extended and optimized by Su et al. [64] to higher dimensions. However, this approach only outperforms DP-Lloyd in datasets with less than three dimensions [64]. Another, more theoretical, line of work focuses on minimizing approximation error bounds, without implementation or empirical evaluation [20, 21, 3, 54, 63, 27]. The work of Ghazi et al. [27] concludes this long line of work by achieving the same approximation factor as the best non-private algorithms. However, these methods suffer from superlinear running times, making them impractical for large datasets, where privacy is most important.
4.2.2 Local DP
@@
In contrast to central DP, the local model operates under the premise that no trusted curator is available. This necessitates that individuals randomize their data locally before it is aggregated, which reduces utility by introducing a significantly higher level of noise (proportional to the number of data points). Several works have been proposed in this area [54, 62, 13], most of which focus on the theoretical error bounds rather than practical applications. Another line of work studies private aggregation (a building block for federated DP $k$-means) in local and shuffle DP models [28, 4]. However, these works cannot overcome the utility gap stemming from the noise being added to each data point. Finally, Li et al. [43] attempt to bridge the gap between these central and local models in the vertically federated setting. Their approach relies on an untrusted curator aggregating noisy local centers and membership encodings from the clients to generate a private synopsis. However, this approach is specifically tailored for the vertical setting and cannot be applied to our configuration. Our work yields utility that is even better than state-of-the-art work in the central model [64] while obtaining the trust assumptions of the local model and operating in the federated setting.
5 Distributed DP $k$-means
@@
First, we overview our modified DP algorithm using the maximum radius constraint and relative updates. Then, we describe FastLloyd, including the initialization, assignment, and update steps. Finally, we provide our security theorem, and discuss how we choose the various parameters.
5.1 Radius Constrained DP $k$-means
In this section, we describe how we improve the clustering utility over the work of Su et al. [64], a contribution that is independently applicable to the central model of DP. We observe that despite being common in the non-private literature [8, 42], neither federated exact $k$-means nor DP $k$-means algorithms in the literature apply any constraints to the $k$-means objective. In this work, we focus on a type of constraint not studied in the non-private literature: bounding the radius of each cluster. This constraint is particularly useful in the context of DP $k$-means, as it allows a tighter sensitivity bound when computing cluster updates relative to the previous centroid.
In the DP $k$-means literature [9, 47, 16, 64], a typical private clustering algorithm is to run Lloyd’s algorithm, with DP noise added to the update step. The DP mean is computed by two DP queries: one for the sum and one for the count. The sum is then divided by the count as a post-processing step. The challenge in these protocols is that the sensitivity of the sum is bounded by the domain size. If the datapoints $x_{l}$ are contained in $[-B,B]^{d}$, then the $l_{2}$ sensitivity of the sum is $B\sqrt{d}$. This is because, with no other constraints, there exists a worst-case data point that could add $B$ to the sum in all dimensions, regardless of the cluster it is assigned to.
In this work, we modify the $k$-means objective such that in addition to minimizing the WCSS (Eqn 1), a constraint on the maximum radius of each cluster must also be satisfied. Specifically, no data point in a cluster can be more than $\eta$ away from the cluster’s centroid. Thus, the objective becomes to minimize (Eqn 1) s.t.:
$||x_{l}-\mu_{j}||_{2}\leq\eta\quad\forall x_{l}\in O_{j},\forall j$ (4)
To enforce this constraint, we modify the assignment step to not assign a data point to a cluster if it is more than $\eta$ away from its nearest centroid. Any unassigned data points are discarded and, therefore, do not contribute to any sum or count query. Intuitively, bounding the maximum radius of each cluster limits how much a cluster can move by adding or removing a data point since a worst-case data point must be close to a cluster’s centroid to be factored into the mean. However, simply bounding the radius of the cluster does not tighten the sensitivity of the sum query. This is because the sensitivity analysis must consider the worst-case cluster. If a cluster is within $\eta$ of the domain boundary, then the $\ell_{2}$ sensitivity of the sum query is still $B\sqrt{d}$.
To realize the reduction in sensitivity from bounding the radius, we must modify the update step as well as the assignment step. Instead of perturbing the sum of the data points themselves, we perturb the difference between the data points and their assigned cluster centroid. Specifically:
$R_{j}^{(t)}=\sum\limits_{x_{l}\in O_{j}^{(t)}}x_{l}-\tilde{\mu}_{j}^{(t-1)}$ (5)
Essentially, we are computing the “updates” to the centroids, rather than the centroids themselves. This simple change has a significant impact on the sensitivity of the sum query. The sensitivity of the relative sum is now bounded by the maximum radius parameter $\eta$:
Theorem 5.1.
If the constraint (4) is satisfied, then:
$\Delta^{R}=\max\limits_{D,D^{\prime}\in\mathcal{D}}||R_{j}^{(t)}(D)-R_{j}^{(t)}(D^{\prime})||_{2}\leq\eta$ (6)
for all clusters $j\in[k]$, and iterations $t\in[T]$.
We prove this in Appendix A.1. The intuition is that a worst-case data point only contributes its distance to the centroid (which is bounded by $\eta$), rather than its distance to the origin.
Our maximum distance constraint naturally bounds the $l_{2}$ sensitivity as it bounds the Euclidean distance, and thus, Gaussian noise is a natural choice that also scales more efficiently to higher dimensions. It also allows for a simple and tight privacy analysis using Gaussian-DP [15]. We use the analysis of Balle and Wang [6, Algorithm 1] to obtain the noise multiplier $\sigma$. We divide the privacy budget between the relative sum and the count following our analysis in Section 5.4 as:
$\sigma^{R}=\frac{\sigma\sqrt{1+\sqrt{4d}}}{\sqrt[4]{4d}}\quad\sigma^{C}=\sigma\sqrt{1+\sqrt{4d}}$ (7)
We then add noise as follows:
$\tilde{R}_{j}^{(t)}=R_{j}^{t}+\gamma^{R}\quad\text{where}\quad\gamma^{R}\sim\mathcal{N}(0,(\sigma^{R})^{2}\eta^{2}T)$ (8)
$\tilde{C}_{j}^{(t)}=C_{j}^{t}+\gamma^{C}\quad\text{where}\quad\gamma^{C}\sim\mathcal{N}(0,(\sigma^{C})^{2}T).$ (9)
We prove in Section 5.3 that adding noise in this way satisfies DP. After adding noise, we compute the new centroid as:
$\tilde{\mu}_{j}^{(t)}=\frac{\tilde{R_{j}}^{(t)}}{\tilde{C_{j}}^{(t)}}+\tilde{\mu}_{j}^{(t-1)}.$ (10)
Without DP noise, this yields an equivalent cluster update.
$\mu_{j}^{(t)}$ $=$ $\frac{S_{j}^{(t)}-C_{j}^{(t)}\mu_{j}^{(t-1)}}{C_{j}^{(t)}}+\mu_{j}^{(t-1)}=\frac{S_{j}^{(t)}}{C_{j}^{(t)}}$ (11)
where $S_{j}^{(t)}=\sum_{x_{l}\in O_{j}^{(t)}}x_{l}$ is the sum of the data points. With the DP noise, we get an additional noise (error) term in the sum compared to noising the sum directly:
$\tilde{S}_{j}^{(t)}$ $=$ $\tilde{R}_{j}^{(t)}+\tilde{C}_{j}^{(t)}\tilde{\mu}_{j}^{(t-1)}$ (12)
$=$ $S_{j}^{(t)}-C_{j}^{(t)}\tilde{\mu}_{j}^{(t-1)}+\gamma^{R}+(C_{j}^{(t)}+\gamma^{C})\tilde{\mu}_{j}^{(t-1)}$ (13)
$=$ $S_{j}^{(t)}+\gamma^{R}+\gamma^{C}\tilde{\mu}_{j}^{(t-1)}.$ (14)
However, as we show in Section 6.2, this additional error is compensated for by the increase in utility from reducing the sensitivity from $B\sqrt{d}$ to $\eta$. The maximum radius further decreases error by reducing the effect of outliers, as data points far away from any centroid will not (and, in some cases, should not) be assigned to any cluster. We also use the maximum radius constraint as an additional post-processing constraint, which we call Radius Clipping. Namely, if a noisy centroid is computed to be more than $\eta$ away from the previous centroid, we truncate it to be $\eta$ away.
5.2 Overview of FastLloyd
In this section, we describe our protocol for federated DP $k$-means, FastLloyd. The focus of FastLloyd is to create the most efficient and accurate protocol possible under the threat model defined in Section 3. Specifically, we refrain from adding additional noise or computations that would be needed to handle client failures or collusion between clients and the server. We leave it to future work to adapt our approach to use more resilient aggregation protocols [10, 39, 61, 22]. Algorithm 1 overviews the protocol from the perspective of a single client. Following the outline of Lloyd’s algorithm, the following sections describe how we design each of its main steps: initialization, assignment, and update.
5.2.1 Initialization: Sphere Packing
@@
We modify the initialization so that it can be carried out in a federated manner. We employ the sphere packing initialization approach of Su et al. [64] as it was shown to outperform random initialization. The sphere packing approach is data-independent and thus does not use any privacy budget. The process can be outlined as follows:
  1. 1.
@@
  3. 3.
    If, after $100$ repeated attempts, it is not possible to find such a point, decrease the radius $a$ and repeat the process.
@@
5.2.2 Assignment: Radius Constrained $k$-means
In Line 7, each client locally computes the closest cluster to each of their data points. If the data point is within $\eta$ of the cluster’s centroid, it is assigned to that cluster (Line 9). Any points further than $\eta$ from the cluster’s centroid are not assigned to any cluster. We discuss how we set $\eta$ in Section 5.5.
@@ 5.2.2 Assignment: Radius Constrained $k$-means
2:Output: Cluster Centres $\tilde{\bm{\mu}}^{(T)}$
[…]
@@
6:     for $x_{l}\in D_{i}$ do
@@
8:         if $||x_{l}-\tilde{\mu}_{j^{\prime}}^{(t-1)}||_{2}<\eta$ then
9:              $O_{ij^{\prime}}^{(t)}\leftarrow x_{l}$
10:     // Local Update
11:     for $j\in\{1,\dots,k\}$ do
12:         Compute $\bar{R}_{ij}^{(t)}=\sum\limits_{x_{l}\in O_{ij}^{(t)}}x_{l}-\tilde{\mu}_{j}^{(t-1)}$
13:         Compute $\bar{C}_{ij}^{(t)}=|O_{ij}^{(t)}|$
14:     // Global Update
15:     $\widehat{\bm{U}}=$ GlobalMSA($\bar{R}_{ij}^{(t)}$, $\sigma^{R}$, seed, $t$)
16:     $\widehat{\bm{C}}=$ GlobalMSA($\bar{C}_{ij}^{(t)}$, $\sigma^{C}$, seed, $t$)
17:     // Post Process Result
18:     for $j\in\{1,\dots,k\}$ do
19:         $\widehat{\mu}_{j}=\frac{\widehat{\bm{U}}}{\widehat{\bm{C}}}+\tilde{\mu}_{j}^{(t-1)}$
20:     $\tilde{\bm{\mu}}^{(t)}=$ Fold($\widehat{\bm{\mu}}$)
5.2.3 Local Update
The client has already locally assigned each of their data points to a cluster, following the constraints in the previous step (Line 9). Next, they compute the relative sum and the count for each cluster using their local dataset. We call this the local update step. The output of the relative sum (Line 12) and the count (Line 13) are two matrices of dimensions $(k\times d)$ and $(k\times 1)$ respectively.
5.2.4 Global Update: Masked Secure Aggregation
We modify the update step to release a perturbed version of relative sums and the counts to each client under the CDP security model. By publishing the previous noisy centroids, the assignment and local update computation can be performed locally by each client instead of using a secure computation protocol. In the global update step (Lines 15 and 16), we privately aggregate (sum) the local updates of all clients, perturb the result, and reveal the noisy global relative sums and counts for the next round.
We describe our aggregation protocol in Figure 1. We call this protocol Masked Secure Aggregation (MSA). In MSA, similar to regular secure aggregation protocols [10], a service provider S securely aggregates values from $M$ clients, all while being oblivious to every client’s contributions. However, in MSA, the server is also oblivious to the result of the aggregation operation; the server only acts as an aggregator (who also adds DP noise) and is oblivious to both the input and output. Each client $p_{i}$ possesses a private value (a matrix $v_{i}\in\mathbf{R}^{k\times d}$, where $\mathbf{R}$ could be $\mathbb{Z}_{2^{64}}$). These clients aim to collectively compute the element-wise sum of their private matrices: $\bm{v}=\sum_{i=1}^{M}{\frac{v_{i}}{M}}$. In addition to the client’s values, we assume the noise scale $\sigma$ and a random $seed$ (that all clients know) are also provided.
Clients Send Data. The first step in the protocol is for each of the $M$ clients to generate a random mask set. The random mask set is composed of $M$ random matrices of the same dimensions as the client items: $\{r_{1},r_{2},\dots r_{M}\}$ where each value is sampled uniformly from $\mathbf{R}$. Because all the clients have access to a shared seed and $PRNG$ they can each compute the entire set locally. In Line 1, each client samples the entire set and sums it (in Line 1) to get the global mask matrix $\bm{r}$ (which will be used to decrypt later).
In Line 1, each private value is converted to a fixed-point format, $\tilde{v_{i}}$, by scaling it up with a power-of-2 scale factor, $2^{q}$. Formally, $\tilde{v_{i}}=\lceil{v_{i}\times 2^{q}}\rfloor$. All fractional values are rounded to the nearest representable value in this fixed-point representation. The choice of the scale factor, $2^{q}$, determines the precision of the representation [^1] A larger $2^{q}$ allows for greater precision but also reduces the number of bits available for the integer part of the number, which might cause overflows. A workaround is to increase the bit-width of the operations, which increases the computational load and the communication cost.. In practice, we empirically choose $q=16$.
Then, in Line 1, each client encrypts (masks) their input $\tilde{v_{i}}$ by adding their share of the mask $r_{i}$. Finally, each client sends their masked values to the server. This step masks the actual value $v_{i}$ from the server because $r_{i}$ acts as a one-time pad. Note that this masked value is never sent to other clients, as they would be able to unmask it easily.
[…]